OpenStudy (anonymous):

If x varies inversely with y and x = 8 when y = 6, find y when x = 10. (Points : 1) a. y=5/12 b. y=4.8 c. y=7.5 d. y=40/3

OpenStudy (paki):

Varying inversely is the opposite of varying directly... so you can solve direct proportions by dividing using: x1/y1 = x2/y2, where x1, y1 are the first set of numbers and x2, y2 is the second set. Likewise, you use multiplication if they vary inversely: x1y1 = x2y2 In this case: (8)(6) = (10)(y) y = 8*6/10 y= 48/10 y= 4.8 @kittykat3

OpenStudy (paki):

but if they had varied directly, you could have solved them like this: 8/6 = 10/y... and would solve it for "y"....